OpenStudy (anonymous):

solve (y+1)(y-3) = -6

OpenStudy (luigi0210):

Welcome to openstudy :) First what you do is FOIL

OpenStudy (anonymous):

foil what part though? I have only had problems that looks like 3(y+1)(y-3)=whatever

OpenStudy (anonymous):

You must expand it out like so: \[(y+1)(y-3)=y(y-3)+1(y-3)=y^2-3y+y-3=y^2-2y+3\] So you have \[y^2-2y+3=-6\] And you can move everything over to one side like so: \[y^2-2y+3+6=0\phantom{spc}\rightarrow\phantom{spc}y^2-2y+9=0\] So now that you have a quadratic equation, you can use the quadratic formula

OpenStudy (anonymous):

I can demonstrate the quadratic equation if you want!

OpenStudy (luigi0210):